OpenStudy (anonymous):

need help with this im lost 1. 1. Find the slope of the line by counting the change in y and the change in x. 2. Find the slope of the line by counting the change in y and the change in x.

OpenStudy (anonymous):

Question 1 chart

OpenStudy (anonymous):

Question 2 chart

OpenStudy (ghostfire):

the equation of the slope is \[\frac{ y_{2} -y _{1}}{x_{2}-x _{1} }\] on the your first graph, the higher point is (x,y)=(-3, 2), the lower point is (-2, -4) so try to plug in the number in the equation and you will get the slope of the line.

OpenStudy (anonymous):

and how do i do that ?

OpenStudy (anonymous):

nvr mind

OpenStudy (anonymous):

so you get -4 - 2/ -2 - (2)

OpenStudy (anonymous):

M= -4 - 2/ -2 - (2) = -6,-4 corrrect right

OpenStudy (ghostfire):

M=(-4-2)/(-2-(-3)) m=(-6)/(-2+3) m=(-6)/(1) m=-6

OpenStudy (ghostfire):

A slope is not a position point, so it won't contain two numbers like (-6,-4). In your first picture the slope is -6, it means for every 1 step you go to the right (positive x), you will go down 6 (negative y).

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what about chart 2